OpenStudy (anonymous):

Show by complete Induction \(\forall n \in \mathbb{N}\): \[\large\sum_{k=1}^{n}kx^{k} = \frac{nx^{n+1}}{x-1}-\frac{x^{n+1}-x}{(x-1)^{2}}\]

OpenStudy (anonymous):

I know i need to start firs with setting n=0 and n=n+1 but i dont get any meaningfull solution

OpenStudy (foolaroundmath):

Try googling for Arithmetico Geometric Progressions

OpenStudy (turingtest):

I think you need k's where you have n's can I edit your Q again?

OpenStudy (anonymous):

yes naturally TuringTest..

OpenStudy (turingtest):

is that right, or were some of those actually supposed to be n's ?

OpenStudy (anonymous):

you are right TuringTest i have one k wrong i edit it wait

OpenStudy (anonymous):

now its ok, if you refresh you have right version..

OpenStudy (turingtest):

am I just messing up or am I getting this as not true for n=1

OpenStudy (anonymous):

I have the complete file in attachment you can take a look

OpenStudy (anonymous):

@TuringTest i understand what u mean right now, i also didnt get a meaningfull solution for n=1 but i am noob at mathematics.. :) maybe i do wrong..

OpenStudy (turingtest):

no it's correct for n=1 I was mistaken

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

@TuringTest can you post your solution pls ?

OpenStudy (turingtest):

I don't have a proof, but you want me to show that it is true for n=1 ?

OpenStudy (foolaroundmath):

Proposition : \(\Large \sum_{k=1}^{n}kx^{k} = \frac{nx^{n+1}}{x-1} - \frac{x^{n+1}-x}{(x-1)^{2}}\) Proposition for n= 1: \(\Large LHS = x, RHS = \frac{1.x^{2}}{x-1}-\frac{x^{2}-x}{(x-1)^{2}}\) \(\Large RHS = \frac{x^{2}}{x-1}-\frac{x(x-1)}{(x-1)^{2}} = \frac{x^{2}-x}{x-1}=x\) Thus the proposition is true for n = 1.

OpenStudy (turingtest):

(I mean I have no complete proof by induction yet)

OpenStudy (turingtest):

yeah, FoolAround beat me to it....

OpenStudy (foolaroundmath):

now, let us assume the proposition is true for \(n=m\). or \[\Large \sum_{k=0}^{m} = \frac{mx^{m+1}}{x-1}- \frac{x^{m+1}-x}{(x-1)^{2}}\] We have to show that the proposition is true for \(n = m+1\) given that the proposition is true for \(n=m\)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

@FoolAroundMath thank you but i couldnt understand how is it possible ?\[\frac{x^{2}-x}{x-1}=x\]

OpenStudy (turingtest):

factor out x on the top

OpenStudy (foolaroundmath):

uggh at the typos: in prev post it should be k=1 to k=m. Rewriting proof now >_<

OpenStudy (turingtest):

you should always copy before posting for that reason ;) you clearly were going in th right direction, stripping out the last term

OpenStudy (foolaroundmath):

and the fact that OS keeps crashing does not help. Writing in notepad and copy pasting now.

OpenStudy (foolaroundmath):

So, we need to show that \[\Large \sum_{k=1}^{m+1}kx^{k} = \frac{(m+1)x^{m+2}}{x-1} - \frac{x^{m+2}-x}{(x-1)^{2}}\] So, lets start from the left hand side of the equation. \[\Large \sum_{k=1}^{m+1}kx^{k} = (m+1)x^{m+1}+\sum_{k=1}^{m}kx^{k}\]Now we know the value of the sum from k=1 to k=m. So substituting that, we get \[\Large \frac{(m+1)x^{m+1}(x-1)}{x-1}+\frac{mx^{m+1}}{x-1} - \frac{x^{m+1}-x}{(x-1)^{2}}\] I've multiplied (x-1) with both numerator and denominator for \((m+1)x^{m+1}\)

OpenStudy (foolaroundmath):

\[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{(m+1)x^{m+1}}{x-1}+\frac{mx^{m+1}}{x-1}-\frac{x^{m+1}-x}{(x-1)^{2}}\] \[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{mx^{m+1}}{x-1}-\frac{x^{m+1}}{x-1}+\frac{mx^{m+1}}{x-1}-\frac{x^{m+1}-x}{(x-1)^{2}}\] \[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{x^{m+1}(x-1)}{(x-1)^{2}}-\frac{x^{m+1}-x}{(x-1)^{2}}\] \[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{x^{m+2}}{(x-1)^{2}}+\frac{x^{m+1}}{(x-1)^{2}}-\frac{x^{m+1}-x}{(x-1)^{2}}\] \[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{x^{m+2}}{(x-1)^{2}}+\frac{x}{(x-1)^{2}}\] \[\Large \frac{(m+1)x^{m+2}}{x-1} - \frac{x^{m+2}-x}{(x-1)^{2}}\]which is what we had to show

OpenStudy (anonymous):

\[\frac{x^{2}-x}{x-1}=x\] how is it possible ?

OpenStudy (foolaroundmath):

\(x^{2}-x = x(x-1) \)

OpenStudy (turingtest):

\[\frac{x^2-x}{x-1}=\frac{x\cancel{(x-1)}}{\cancel{x-1}}=x\]

OpenStudy (anonymous):

aah ok thanks @FoolAroundMath and @TuringTest

OpenStudy (foolaroundmath):

@TuringTest what is the LaTeX for strikeout ?

OpenStudy (anonymous):

@FoolAroundMath i think you mean this, its with cancel \frac{x\cancel{(x-1)}}{\cancel{x-1}}=x